A battery powered torch uses a 9 V battery and dissipates 0.25 W of electrical power in the bulb.

If the torch is used for 1.0 hours then how much charge has passed through the bulb?



1
Expert's answer
2022-05-26T14:07:24-0400

The electrical current through the bulb

"I=\frac{P}{V}=\frac{0.25}{9}=0.0278\:\rm A"

The charge has passed through the bulb

"q=It=0.0278*3600=100\:\rm C"